Project

General

Profile

10/08/2014

(Notes Converted from handwriting)
Attending: Mat, Louise, Keith, Jose, Gavin and Fernanda from Roc
-Kyle: reports progress, but nothing
to report today. Update at next DQ Meeting.
Far Detector
-10/2-10/3: FEB hit rate, hits per spill low
- DB 14 failing good run selection
- Gain study

Gain Study

HW Watch: consistent set of bad channels compared to last week

Near Det- DCM 4/1+ cooling off over weekend
10/6 1 DCM off over weekend, back on 10/6
→Number of hits per spill high after

AOB
Louise- Working with Fernanda to make ND swap list.
-Beam back on 10/21 at soonest

Fernanda needs to setup to processing for activity triggers for ND
•will manually produce maintenance list
working to getting DCM sync plots with Jose.
Site wide power outage on Friday.
Keith- storing noise info...

Wednesday, October 8, 2014
12:25 PM